OpenStudy (anonymous):

How would you go about solving this: p(a + 1) = a^2 +2a +1 What is p(x)?

jhonyy9 (jhonyy9):

a^2 +2a +1 = ? like a perfect square

jhonyy9 (jhonyy9):

(a+1)^2 =

OpenStudy (anonymous):

hmm, maybe I'm misunderstanding you but p is a function

OpenStudy (anonymous):

so if the function p(a + 1 ) = a^2 +2a +1, what would p(x) give? I may even be misunderstanding the entire question...

OpenStudy (anonymous):

jhonyy9 on the right track

OpenStudy (anonymous):

is*

OpenStudy (anonymous):

a better way to look at your problem is p(a+1)=(a+1)^2